Cercarial dermatitis, commonly known as swimmers itch, is a rash caused by an allergic reaction to parasites that burrow i
Answer #1

\widehat p = 44/127 = 0.3465

For 94% Confidence interval, the critical value is z* = 1.88

The upper limit of the 94% confidence interval is

\widehat p + z^{*} * \sqrt{\frac{\widehat p(1 - \widehat p)}{n}}

= 0.3465+ 1.88* \sqrt{\frac{0.3465(1 - 0.3465)}{127}}

= 0.3465 + 0.0794

= 0.4259